# Quantum-tunnelling deep neural networks for sociophysical neuromorphic AI

### v1

@@ Abstract
The discovery of the quantum tunnelling effect—the transmission of particles through a high potential barrier—was one of the most impressive achievements of quantum mechanics made in the 1920s. Responding to the contemporary challenges, I introduce a novel deep neural network (DNN) architecture that processes information using the effect of quantum tunnelling. I demonstrate the ability of the quantum tunnelling DNN (QT-DNN) to recognise optical illusions like a human. Hardware implementation of QT-DNN is expected to result in an inexpensive and energy-efficient neuromorphic chip suitable for applications in autonomous vehicles. The optical illusions recognition tests developed in this paper should lay foundations for cognitive benchmarking tasks for AI systems of the future, benefiting the fields of sociophysics and behavioural science.
@@ I.1 Can AI see optical illusions?
How will a conscious robot [1] see us and our world? Will it understand how humans behave in natural complex environments such as oceans and space [2, 3, 4] and make decisions in virtual reality systems [5, 6, 7] and video games [8]? Speaking broadly, will a robot differentiate lie from truth in media, social networks and politics [9, 10, 11]?
Biological vision is enabled by physical, physiological and psychological processes [12]. In turn, AI employs machine vision [13] and neural network models [14] to recognise and classify objects. Hence, even though machine vision shares certain features with biological vision [15], in general AI and humans see two very different worlds [16, 17, 18].
For example, how would AI see the Necker cube [19, 20] and Rubin’s vase [21, 22] (Fig. 1a, b)? While a human looking at these paradigmatic optical illusions [19] can easily report a random switching of their perception from one possible interpretation of the figure to another [20, 23] (Fig. 1c), AI will struggle to produce a similar output in spite of the advances in simulation of human vision [24, 25, 26, 27, 28, 29, 30].
Figure 1: (a) The Necker cube: The answer to the question ‘Is the shaded face of the cube at the front or at the rear?’ will randomly switch between two stable perceptual states corresponding to the front ($|0\rangle$) and rear ($|1\rangle$) face of the cube. (b) Rubin’s vase: Do you see two people looking towards each other ($|0\rangle$) or a vase ($|1\rangle$)?’ (c) According to the traditional psychological theories, the change from one perceptual state to another is instantaneous (the dotted line). However, current research works demonstrate that humans might see a superposition of the states $|0\rangle$ and $|1\rangle$ (the solid curve).
@@ I.2 AI, psychology and quantum mechanics
AI cannot see optical illusions since its algorithms do not take into account peculiar psychological and neurological aspects of human vision [31, 19, 32, 22]. Yet, although this problem has been stressed in the literature [33, 34, 35, 36], the relevance of visual and biological complexity of natural environments to AI has not been fully appreciated [37, 38, 4].
Experimental studies [39, 40, 41] demonstrated that the perception of the Necker cube and other ambiguous figures may not abruptly switch between the states $|0\rangle$ and $|1\rangle$ (the dotted line in Fig. 1c) but continuously oscillates between them (the solid curve in Fig. 1c). Those results were interpreted as the ability of humans to see a quantum-like superposition of $|0\rangle$ and $|1\rangle$ states [42, 43, 20, 44]. Indeed, in psychological experiments humans are asked to push an electric button every time their perception of the Necker cube switches between $|0\rangle$ and $|1\rangle$. While humans might see the cube in a superposition state [31, 41], by pressing the button the observer ‘collapses’ the superposition to one of the possible ‘classical’ states [23]. This means that the ability of humans to access and use environmental information is limited, which may result in ambiguous interpretation. Subsequently, our perceptual system needs to disambiguate and interpret the available sensory information to construct stable and reliable percepts [45]. Yet, this suggests that AI can help humans to enhance their perception and gain access to information that our senses and brain cannot process naturally.
Further attempts to understand bistable perception using the methods of quantum mechanics [43, 20, 44] resulted in the suggestion to study the superposition of the perceptual states using a quantum oscillator model (QOM) [20]. Compared with Markov models [20], QOM accounts for multiple outcomes while processing input data with a large set of constraints, thus describing human mental states [43, 44] more efficiently than classical models.
The predictions of QOM are in good agreement with a model based on the quantum Zeno effect [42]. Besides, the ability of QOM to simulate the perception of optical illusions was improved [46, 23] by combining harmonic motion with the effect of quantum tunnelling (QT) through a potential barrier [47, 48]. Since both QOM and quantum Zeno effect underpin quantum neural network models [49], the predictions made by QOM were additionally validated using a deep neural network (DNN) model [23] combined with a quantum generator of truly random numbers [50].
@@ I.3 What is this article about?
In this work, I suggest and theoretically validate a novel neuromorphic DNN model (see Sect. II.1 for state-of-the-art) that processes information using the physical effect of QT. I showcase the ability of the QT-DNN model to enable AI to recognise practically important optical illusions, the perception of which should pave the way for the development of machine vision system capable of recognising more complex visual effects used in fine arts and cinematography. I also demonstrate that QT-DNN can be implemented as a hardware neuromorphic chip.
My findings can be used in AI system and brain-computer interfaces aimed to enhance the performance of astronauts in long-duration spaceflight [3] and help airline pilots be aware of disorientation, loss of perspective and misinterpretation of data provided by flight instruments [51]. Interestingly enough, the algorithms proposed in this work can also help conduct research on gender since it has been demonstrated that fluidity of gender identity might be induced by optical illusions [52].
Finally, concluding the paper with a discussion of the relevance of the effect of QT to the hypothesised brain-mind link [53, 54], I contribute to bridging the gap between the theories of quantum mind and decision-making [43, 20] and the practically important field of sociophysics [9, 11].
@@ II QT-DNN architecture
II.1 Neuromorphic quantum DNN models
DNNs—networks composed of several layers (Fig. 2a) [14]—are pivotal for machine vision and human-computer interactions [13]. Yet, they are mathematically complex [55], require considerable computational resources [56] and consume significant electric power [57].
The efficiency of DNNs can be increased by constructing the network following the principles of quantum computing [58, 59, 60] and optimising their algorithms using the fundamental laws of physics [61]. DNNs can also be built using the principles of neuromorphic computing [62, 63].
Neuromorphic computers mimic the operation of a biological brain [62, 64]. Several neuromorphic DNN models have been demonstrated [65, 66], where, instead of relying on conventional Boolean arithmetic, information is processed using nonlinear dynamical properties of a physical system [64, 63]. Such DNN models can be implemented as an inexpensive and low power consuming chip [65, 67] suitable for applications in AI and robotics [66].
II.2 Quantum tunnelling effect
In quantum mechanics [48], the rectangular potential barrier is a computational problem that demonstrates the phenomenon of QT (Fig. 2b). The problem consists in solving a one-dimensional time-independent Schrödinger equation for an electron that encounters a rectangular potential energy barrier [47]

### v2

# Quantum-tunnelling deep neural network for optical illusion recognition

@@ Abstract
The discovery of the quantum tunnelling (QT) effect—the transmission of particles through a high potential barrier—was one of the most impressive achievements of quantum mechanics made in the 1920s. Responding to the contemporary challenges, I introduce a deep neural network (DNN) architecture that processes information using the effect of QT. I demonstrate the ability of QT-DNN to recognise optical illusions like a human. Tasking QT-DNN to simulate human perception of the Necker cube and Rubin’s vase, I provide arguments in favour of the superiority of QT-based activation functions over the activation functions optimised for modern applications in machine vision, also showing that, at the fundamental level, QT-DNN is closely related to biology-inspired DNNs and models based on the principles of quantum information processing.
@@ I.1 Can AI see optical illusions?
How will a conscious robot [1, 2] see us and our world? Will it understand how humans behave in natural complex environments such as oceans and space [3, 4, 5] and make decisions in virtual reality systems [6, 7, 8] and video games [9]? Speaking broadly, will a robot differentiate lie from truth in media, social networks and politics [10, 11, 12]?
Biological vision is enabled by physical, physiological and psychological processes [13]. In turn, AI employs machine vision [14] and neural network models [15] to recognise and classify objects. Hence, even though machine vision shares certain features with biological vision [16], in general AI and humans see two very different worlds [17, 18, 19, 20, 21].
For example, how would AI see the Necker cube [22, 23] and Rubin’s vase [24, 25] (Fig. 1a, b)? While a human looking at these paradigmatic optical illusions [22] is able to report a random switching of their perception from one possible interpretation of the figure to another [23, 26] (Fig. 1c), AI will struggle to reproduce the human perception in spite of the advances in simulation of human vision [27, 28, 29, 30, 31, 32, 33].
Figure 1: (a) The Necker cube: The answer to the question ‘Is the shaded face of the cube at the front or at the rear?’ will randomly switch between two stable perceptual states corresponding to the front ($|0\rangle$) and rear ($|1\rangle$) face of the cube. (b) Rubin’s vase: Do you see two people looking towards each other ($|0\rangle$) or a vase ($|1\rangle$)?’ (c) As per the traditional theory, the change from one perceptual state to another is binary (the dotted line), i.e. it is from $|0\rangle$ to $|1\rangle$ and vice versa. However, the current research works demonstrate that humans might see a superposition of the states $|0\rangle$ and $|1\rangle$ (the solid curve).
@@ I.2 AI, psychology and quantum mechanics
AI cannot see optical illusions like a human since its algorithms do not take into account psychological and neurological aspects of human vision [34, 22, 35, 25]. An increasing number of research works have also suggested that this discrepancy is likely to arise due to the fundamental differences between artificial and biological neurons [18, 36, 37, 38, 39, 40]. Although this problem has gained more attention in the recent years [41, 42, 43, 44], the impact of visual and biological complexity of natural environments on the ability of AI to see like a human has not been fully appreciated [45, 46, 5].
Experimental studies [47, 48, 49, 50, 51] demonstrate that the perception of the Necker cube and other ambiguous figures may not directly switch between the states $|0\rangle$ and $|1\rangle$ (the dotted line in Fig. 1c) but can continuously oscillate between them (the solid curve in Fig. 1c). Those results have been interpreted as the ability of humans to see a quantum-like superposition of $|0\rangle$ and $|1\rangle$ states [52, 53, 23, 54]. Indeed, in psychological experiments humans are asked to push an electric button every time their perception of the Necker cube switches between $|0\rangle$ and $|1\rangle$. While humans might see the cube in a superposition state [34, 51], by pressing the button the observer ‘collapses’ the superposition to one of the possible ‘classical’ states [26]. This means that the ability of humans to access and process environmental information is limited, which may result in ambiguous interpretation. Subsequently, our perceptual system needs to disambiguate and interpret the available sensory information to construct stable and reliable percepts [55]. Yet, this suggests that AI can help humans enhance their perception and gain access to information that our senses and brain cannot process naturally. Such AI systems, including augmented cognition systems [56, 57] and decision-making support systems [58], can be used to make high-quality administrative, political, financial and military decisions, taking in to account difficult to foresee scenarios of social and cultural interactions, geopolitical landscapes and battlefields in foreign lands [57].
Further attempts to understand bistable perception using the methods of quantum mechanics [53, 23, 54] resulted in the suggestion to study the superposition of the perceptual states using a quantum oscillator model (QOM) [23]. Compared with Markov models [23], QOM accounts for multiple outcomes while processing input data with a large set of constraints, thus describing human mental states [53, 54] more efficiently than the classical models.
The predictions of QOM are in good agreement with the results obtained using the models based on the quantum Zeno effect [52]. Besides, the ability of QOM to simulate the perception of optical illusions was improved [59, 26] by combining harmonic motion with the effect of quantum tunnelling (QT) through a potential barrier [60, 61]. Since both QOM and quantum Zeno effect underpin quantum neural network models [62], the predictions made by QOM were additionally validated using a deep neural network (DNN) model [15, 26] combined with a quantum generator of truly random numbers [63].
@@ I.3 What is this article about?
In this work, I suggest and theoretically validate a novel neuromorphic [64, 65] DNN model that processes information using the physical effect of QT. I showcase the ability of the QT-DNN model to enable AI to recognise practically important optical illusions, the perception of which should pave the way for the development of machine vision systems capable of recognising more complex visual effects used in fine arts and cinematography.
My findings can be used in AI systems and brain-computer interfaces aimed to enhance the performance of astronauts in long-duration spaceflight [4] and help airline pilots be aware of disorientation, loss of perspective and misinterpretation of data provided by flight instruments [66]. Since the analysis of ambiguous figures helps identify people who suffer from mild cognitive impairment, the algorithm proposed in this work can be employed in AI systems used by health professionals to diagnose debilitating diseases such as dementia [67]. Interestingly enough, it can also help conduct research on gender since it has been demonstrated that fluidity of gender identity might be induced by optical illusions [68].
@@ II QT-DNN architecture
II.1 QT-DNN algorithm
Unless otherwise specified, the QT-DNN network used in this paper (Fig. 2a) consists of an input layer that has $L=100$ input nodes, three hidden layers each of which has $N=20$ nodes and an output layer that has $M=2$ output nodes that are used to classify the input dataset. The weights of the connections of the network are updated using a back-propagation training algorithm [15].
The activation function $\phi_{QT}$ of the nodes of the hidden layers is given by the algebraic expressions for the transmission coefficient $T$ of an electron penetrating a potential barrier (Fig. 2b). These expressions are well-known [60, 69] and they are presented in the Appendix section for the sake of self-consistency of the current paper.
The output nodes of QT-DNN are governed by the Softmax function [15]
$\phi_{smax}(v_{i})=\frac{\exp(v_{i})}{\sum_{k=1}^{M}\exp(v_{k})}\,,$ (1)
The network is trained and exploited as follows. First, I construct the output nodes that correspond to the correct answers to the training datasets. Then, I initialise the weights of the neural network in the range from –1 to 1 using a random number generator. Entering the input data $x_{j}$ and the corresponding training data points $d_{i}$, I calculate the error $e_{i}$ between the output $y_{i}$ and target $d_{i}$ as $e_{i}=d_{i}-y_{i}$. Then, propagating the output $\delta_{i}=e_{i}$ in the backward direction of the network, I compute the respective parameters $\delta_{i}^{(n)}$ of the hidden nodes using the equations $e_{i}^{(n)}=W^{{(n)}^{\top}}\delta_{i}$ and $\delta_{i}^{(n)}=\phi_{QT}^{\prime}\left(v_{i}^{(n)}\right)e_{i}^{(n)}$, where the index $n$ denotes the sequential number of the hidden layer, prime denotes the derivative of the activation function and $W^{\top}$ is the transpose of the matrix of weights corresponding to each relevant layer of the network. I continue the back-propagation process until the algorithm reaches the first hidden layer and then I update the weights using the learning rule $w_{ij}^{(n)}\leftarrow w_{ij}^{(n)}+\Delta w_{ij}^{(n)}$, where $w_{ij}^{(n)}$ are the weights between an output node $i$ and input node $j$ of the $n$th layer and $\Delta w_{ij}^{(n)}=\alpha\delta_{i}^{(n)}x_{j}$. These computational steps are sequentially applied to all training data points.
II.2 Neuromorphic QT-DNN model of perception
Based on these facts, it was demonstrated that the introduction of chaotic changes in the architecture of a neural network enables modelling the dynamics of information perception [27, 26]. Thus, I randomly generate the weights $W^{(n)}$, then I train QT-DNN on the $10\times 10$ pixel unambiguous images of the Necker cube and then I exploit it to predict the perceptual state of the ambiguous Necker cube. This procedure is repeated in a loop to plot the dynamics of the perceived states (Figure 3).
To create a more realistic model of quantum dynamics, I employ a physical generator of random numbers based on coherent quantum-optical processes [63] (the emission of coherent light can be associated with the effect of QT [81]). Unlike a pseudo-random generator, a quantum generator produces truly random numbers [63]. This ensures that the software model of QT-DNN is not biased towards one of the possible perceptual states of optical illusions [27, 82].
III Results
Figure 4a plots the simulated probability of perceiving $|0\rangle$ and $|1\rangle$ states of the Necker cube as a function of time (arbitrary time units are used in this work; for a relevant discussion of the cognitive timescale see Refs. [52, 83]). The nondimensionalised thickness of the potential barrier employed as the activation function of QT-DNN is $\sqrt{2mV_{0}}a/\hbar=$ 0.5 (see the Appendix section for details; for a relevant discussion of the role of the barrier parameters in the quantum models of cognition see Refs. [26, 11]).
The probability curves were obtained as a result of 40 consecutive computational runs of the algorithm outlined in Fig. 3. The states of the two output nodes of QT-DNN were recorded at the end of each run. Every pair of those data points was computed using a unique set of neural weights $W^{(n)}$. The same procedure was followed to simulate the perception of Rubin’s vase (Fig. 5). The same values of $W^{(n)}$ were used in Fig. 4 and Fig. 5.
In Fig. 4a, we can observe a time-dependent switching between the fundamental perceptual states $|0\rangle$ and $|1\rangle$ of the Necker cube. In agreement with the predictions of the quantum cognition theory [23], the switching between $|0\rangle$ and $|1\rangle$ is not binary (i.e. it does not occur directly from $|0\rangle$ to $|1\rangle$ and vice versa) but it involves intermediate states that can be interpreted as a superposition of $|0\rangle$ and $|1\rangle$. At some instances of time (e.g., $T=17$), the model predicts an important, from the point of view of the quantum cognition theory, result where the probability of simultaneously perceiving the two possible states of the cube oscillates around 0.5 [23, 26].
The results produced by QT-DNN can be used to simulate the classical perception of the Necker cube, i.e. the perception that admits only two binary states of the cube [26]. To that end, it can be assumed that the classical perception Percept = 1 when the quantum probability $P_{|1\rangle}\geq 0.5$ and Percept = 0 when $P_{|1\rangle}<0.5$.
Figure 4: (a) Perceptual switching data (with the lines serving as a guide to the eye) produced by QT-DNN trained on the Necker cube. The data points located in between 0 and 1 are in a superposition of the states $|0\rangle$ and $|1\rangle$. (b) Classical perception computed using the procedure explained in the main text.
In Fig. 4b, we can see that the classical perception curve consists of a series of square-like pulses of varying duration. This result should plausibly describe the human perception of the Necker cube. However, the classical result does not contain important information that can be used to better understand how humans control their perception of the illusion. In fact, the perception of the Necker cube is subjective and it may also be affected by the previous experience of the observer with ambiguous drawing and observer’s attempts to force the switching to occur by focusing on different parts of the cube [34, 22] and voluntarily blinking [75].
As can be seen in Fig. 5, the pattern of the perception switching of Rubin’s vase is similar to that of the Necker cube. However, QT-DNN predicts a higher switching frequency and longer periods of time corresponding to a superposition of the $|0\rangle$ and $|1\rangle$. This result is consistent with the theory that states that Rubin’s vase combines an optical illusions with the perception of background [25]. Indeed, observers may interpret this figure not only as faces-vase but also as a white (black) vase on black (white) background, which gives them an additional point of reference that can be used to better disambiguate the drawing [84, 85].
Figure 5: The same as in Fig. 4 but for Rubin’s vase.
IV Discussion
In this section, I demonstrate that, in the context of cognitive and decision-making DNN models, a QT-based activation function provides advantages over the standard activation functions employed in the field of machine learning [15]. I also show that, at the fundamental level, QT-DNN operates similarly to theoretical models of biological neural networks [86, 38]. Although a solid link between the theory of quantum cognition and neuroscience is yet to be established, an extended discussion surveying the recent advances in the adjacent areas of quantum information and quantum consciousness (Sect. V) provides arguments in favour of its existence.
Figure 6: Probability of perceiving the $|1\rangle$ state of (a) the Necker cube and (b) Rubin’s vase computed using DNN software that employs QT, ReLU and Sigmoid activation fucntions. (c) Similarity between the perception switching curves calculated using the DTW algorithm.
Several DNN models of human perception of optical illusions were proposed in the previous works [27, 28, 41, 29, 30, 31, 32, 33, 26]. Each model produced feasible results. However, no quantitative assessment of the accuracy of those models against psycho-physiological datasets has been carried out due to the unavailability of reliable human-generated data. Subsequently, as discussed above, experimental brain activity data have been used to evaluate the performance of the models. Thus, even though in its traditional formulation the quantum cognition theory does not imply any direct association with the function of a biological brain [23, 54], it is conceivable that quantum cognition and neuroscience should be inter-related, at least in some aspects [87, 88, 89].
Many DNN models adopted in the studies of optical illusions have been inspired by the advances made in the field of machine vision, thereby employing the advanced activation functions such as ReLU [41, 33, 26]. However, it has been recognised that the application of ReLU and other standard activation functions in DNN models of cognition and brain activity may lead to incorrect results [38, 39, 40]. Yet, although Sigmoid-like activation function are also employed in DNN models [15], their application in the studies of optical illusions appears to be motivated not by the knowledge from the field of machine learning but by the fact that similar functions have been used in certain models of biological neurons [86, 38].
To provide a unified comparison, in Fig. 6a, b I plot the probability of perceiving the $|1\rangle$ state of the Necker cube and Rubin’s vase, respectively, computed by the DNN models that employ QT, ReLU and Sigmoid activation functions (all DNNs were trained under otherwise identical conditions). To quantify the similarity between the calculated switching curves, I employ a dynamic time warping (DTW) time-series analysis algorithm [90] that enables measuring similarity (distance) between temporal sequences that vary in speed [91]. The choice of DTW is justified because this algorithm has often been applied to human-generated time series such as speech audio signals and walking dynamics [91]. Thus, as shown in Fig. 6c, the most similar results were produced by QT-DNN and Sigmoid-DNN but the largest distance was observed between the outputs of QT-DNN and ReLU-DNN.
The perception switching patterns produced by ReLU-DNN resembles the classical binary perception patterns shown in Fig. 4c and Fig. 5c: the output of ReLU-DNN is predominantly a pure $|0\rangle$ or $|1\rangle$ state, with just a few data points corresponding to a superposition of $|0\rangle$ and $|1\rangle$. Thus, even though ReLU-DNN might have demonstrated an optimal result in terms of the performance metrics adopted in the field of machine learning, its output disagrees with the predictions a large and growing body of quantum models of perception. Moreover, in line with other relevant works [18, 36, 37], this discrepancy may indicate that the use of ReLU in DNN models designed to recognise optical illusions is unjustified.
On the other hand, a similarity of the outputs produced by QT-DNN and Sigmoid-DNN suggests that the effect of QT serves as a more adequate activation function for the analysis of psycho-physiological data. The following discussion provides additional arguments in support of this statement.
Firstly, as shown in Fig. 7, the shape of the QT activation function is similar to the shape of the neural activation function motivated by biological data [92, 38]. The functionality behind such a shape enables the model to satisfactorily account for the relation between the firing rate of a biological neuron and the input current, also considering the minimal time between two action potentials and other biological mechanisms [86, 38].
Secondly, it has been demonstrated that neural networks based on a Sigmoid-like activation function can simulate such important operations of the quantum decision-making theory as conjunction fallacies (violation of the laws of the classical probability theory by human decision-making) [88]. While the authors of the cited paper might choose a Sigmoid-like activation function due its application in the models of biological neurons, their choice is also consistent with both regular techniques of probability estimation in social-economic research (e.g., with logistic regression [93]) as well as with novel theoretical approaches that analyse human behaviour through the prism of quantum physics [94].
Admittedly, the comparative analysis conducted above does not eliminate the need for a detailed benchmarking of the model outputs against rigorous experimental data. It is likely that such a comparison will be made in the future. Nevertheless, especially because further examples of the application of Sigmoid-like functions [95] and the effect of QT [59, 26] in quantum models of decision-making can be presented, it possible to claim that the results presented in this paper help establish a link between deep learning techniques and quantum cognition theories, also making us one step closer to the development of conscious AI, the potential benefits of which have been anticipated in the Introduction section.
V Conclusions and Outlook
In this paper, I pointed out the possibility of employing the effect of quantum tunnelling as an activation function of artificial neural networks. I demonstrated that the so-constructed neural networks enable accurate modelling of human perception of optical illusions.
During the peer-review process, my attention was brought to the fact that a digital computer running software that implements the QT-DNN algorithm technically remains a classical system, even though the model is based on the solution of a Schrödinger equation and employs random number sets generated by a purely quantum physical system. The classical nature of QT-DNN software is inconsequential in the frameworks of sociophysics and quantum cognition theory, where, generally speaking, a model remains valid as long as the idealisations made in it and the assumptions about the data of interest hold true [11]. Indeed, it has been shown that empirical psychological data obtained in experiments aimed to reveal imprecision of human preference can be robustly explained using a classical electrodynamical model of an electron device based on the purely quantum phenomenon of spin-transfer torque [94].
However, the distinction between classical and quantum systems becomes more prominent in the framework of the theories [69, 96, 97, 98, 99, 100, 101, 102, 103] attempting to explain human consciousness and brain function from the point of view of the quantum information theory [104]. Although the quantum cognition theory and the quantum brain theory are formally independent one of another [23, 54], they should converge at a certain point, resulting in a unified theory [87, 88, 105]. While the task of such a unification would be too ambitious for this paper, also going well beyond its scope, I note that a potential implementation of QT-DNN in hardware should result in a purely quantum neuromorphic computing system based on an analogue platform. Although a quantum neuromorphic system is different from a conventional quantum computing architecture [106], at the conceptual level its hardware implementation should comply with the definition of ‘genuine quantum hardware’ [104] identified in the framework of the quantum brain theory as an essential criterion for understanding consciousness [96].
Data Availability Statement
The computational code that implements QT-DNN and relevant data can be found at https://github.com/IvanMaksymov/DeepTunnel (accessed on 26 July 2024)
Appendix A Quantum tunnelling as an activation function
The electron tunnelling behaviour can be quantified by finding the transmission coefficient from the solution of a time-independent Schrödinger equation